. Problem 2.22 Suppose that a(t) 0.1t? + 1. At time 0, $1,000 is invested. An additional investment of SX is made at time 6. If the total accumulated value of these two investments at time 8 is $18,000, find X ntcrest
Answer #1

Initial Investment = $ 1000, a(t) = 0.1 x (t)^(2) + 1 and Tenure of Investment = t = 8 years

Final Value of Initial Investment when t = 8 will be FV(8) = 1000 x [0.1 x (8)^(2) + 1] = $ 7400

An amount $ X is invested at t=6 and the total value of both investments is $ 18000 at t=8, of which $ 7400 is from the initial investment. Remaining is the final value of investment X

Therefore, FV(X) = 18000 - 7400 = $ 10600

The investment X is made for a duration of 2 years (8-6 = 2) and hence the appropriate value of t in a(t) is 2

Therefore, X x [0.1 x (2)^(2) + 1] = 10600

X = 10600 / 1.4 ~ $ 7571.43